OpenStudy (abbles):

Will medal!! Cylinder A has a radius of 12 inches and a height of 6 inches. Cylinder B has a volume of 648π. What is the percent change in volume between cylinders A and B?

OpenStudy (abbles):

a. Cylinder B is 50% smaller than cylinder A. b. Cylinder B is 75% smaller than cylinder A. c. Cylinder B is 150% bigger than cylinder A. d. Cylinder B is 200% bigger than cylinder A.

OpenStudy (abbles):

Cylinder B is smaller, so I'm thinking it's either a or b?

OpenStudy (anonymous):

The answer is B, because half of A is no where near B

OpenStudy (abbles):

But are you sure B is the right answer? It doesn't seem right to me...

OpenStudy (campbell_st):

well find the volume of the cylinder...

OpenStudy (anonymous):

I'm positive, A is 800pi and B is 600 but B is not 50% of A for sure, which narrows it down to 75%

OpenStudy (abbles):

100.48 is the volume of cylinder A... so what would I do from there?

OpenStudy (sidsiddhartha):

volume of a cylinder\[\pi*r ^{2}*h\]

OpenStudy (anonymous):

12^2 * 6

OpenStudy (anonymous):

clearly not 100

OpenStudy (anonymous):

then ignore the pi and just put it next to it

OpenStudy (abbles):

My mistake... 864pi sorry

OpenStudy (sidsiddhartha):

now find the percentage change of volume

OpenStudy (abbles):

what is the formula for that?

OpenStudy (anonymous):

Step 1: Divide the New Value by the Old Value (you will get a decimal number) Step 2: Convert that to a percentage (by multiplying by 100 and adding a "%" sign) Step 3: Subtract 100% from that

OpenStudy (sidsiddhartha):

%change in volume= \[\frac{ VA-VB }{ VA }*100\]

OpenStudy (anonymous):

You really don't need to find the percentage of change because it becomes quite obvious that it't not 50%

OpenStudy (abbles):

648/864 = .75 .75 x 100 = 75% So the answer is b then? @sidsiddhartha

OpenStudy (anonymous):

Yes B is the answer

OpenStudy (sidsiddhartha):

v=864-648/864=25% so Cylinder B is(100-25)%= 75% smaller than cylinder A

OpenStudy (abbles):

thanks guys

OpenStudy (sidsiddhartha):

^_^